OpenStudy (studygurl14):

I DON'T KNOW HOW TO DO THIS! MEDAL! @jigglypuff314 @dan815 @jim_thompson5910

OpenStudy (studygurl14):

Is there a formula for this, because I really do not want to draw a tree.

OpenStudy (studygurl14):

All I can figure out is that there are 216 total possible outcomes, and 1/6 chance of getting a six, and 5/6 of not getting a six...but the whole thing about "if the first number cube shows 6" is throwing me off

OpenStudy (studygurl14):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

We are told that one cube already shows 6 Since we want exactly 2 cubes to show 6, this means things get boiled down to this question: "what is the probability of exactly one cube showing six?"

jimthompson5910 (jim_thompson5910):

ignore that 3rd cube we have 2 dice probability of rolling a 6 = 1/6 probability of not rolling a 6 = 5/6 probability of rolling 2 dice to get exactly one six = (1/6)*(5/6) = 5/36

jimthompson5910 (jim_thompson5910):

sorry i miscounted, it's better to use a chart like this